OpenStudy (anonymous):

how do you use the quadratic formula to solve 3x^2+4x=6

OpenStudy (anonymous):

move the 6 to the other side so you have \[3x^2+4x-6=0\] then you have your A,B, and C

OpenStudy (anonymous):

how do you finish the problem

OpenStudy (anonymous):

\[ax^2 +bx +c\]

OpenStudy (anonymous):

"-b +/- sqrt(b^2-4ac)" all divided by "2a"

OpenStudy (anonymous):

Set it equal to 0 3x^2 +4x -6 =0 Now use the quadratic formal to solve for x. x= (-4+-sqroot[4^2-4(3)(-6)]/6 =(-4+-sqroot[16+72]/6 =(-4+-sqroot[88]/6: x=(-4+sqroot{88]/6 or x=(-4-sqroot[88]/6 simplify radicals. sqroot[88] =sqroot[2^3x11] = 2sqroot[2x11]= 2sqroot[11]: x=(-4+ 2sqr[22])/6 or x=(-4-2sqr[22])/6 x=((1/3)sqrt[[22]-2]) or x=((1/3) (-2-sqrt[22]))